# A simple model to investigate jet quenching and correlated errors for centrality-dependent nuclear-modification factors in relativistic heavy-ion collisions

### v1

@@ I Introduction
In this work we focus primarily on the centrality and $p_{T}$ dependence of the jet nuclear-modification factor, $R_{AA}$, measured for Pb+Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV, as measured by the ATLAS Collaboration using the anti-$k_{T}$ algorithm with jet radius, $R=0.4$ [23]. These data were selected for the wide range in centrality and $p_{T}$ coverage and quantitative assessment of each of the systematic error components. Final fit results will also be compared to similar measurements by the CMS and ALICE collaborations [24, 25] collaborations. The simple model employed is an extension of the approach adopted in [20], in which the jet-energy loss distribution is described by a $p_{T}$ dependent function and applied to p-p jet cross-section to fit central $R_{AA}$ values using Bayesian methods. This model is then modified to incorporate geometric information provided by the TRENTo  model [16] and the parameters are refit to the centrality-dependent $R_{AA}$ values with full systematic errors treated in two different ways. The jet-quenching distributions of the model are described in Sec. II, and Sec. III provides comparisons to central $R_{AA}$ measurements. Centrality dependent extensions and comparisons are provided in Sec. III, and the summary and conclusions are given in Sec. V. The treatment of systematic error correlations and impacts on fits are explored further in the Appendices.
@@ II Jet-quenching Distributions
The first step in developing a simple model for energy loss is to parameterize unquenched, inclusive jet distributions in $\sqrt{s_{NN}}=5.02$ TeV proton-proton collisions. For this we use the ATLAS measurements in the rapidity range $|y|<2.1$ [26], and fit to Eq. 1 for values of $800100<p_{T}<$ 800 GeV, sufficient to cover the ATLAS jet-$R_{AA}$ measurements spanning $631158<p_{T}<$ 631 GeV,
$\frac{dN^{2}_{\rm jet}}{dp_{T}dy}=\frac{(c_{1}/p_{T})^{[c_{2}+c_{3}\log(p_{T})]}}{1+\exp[(p_{T}-c_{4})/c_{5}]}.$ (1)
@@
The gamma-distribution is motivated by its equivalence to a convolution of $k$ exponentials, thereby providing a quasi-physical interpretation as the number of parton scatterings. There is no physical interpretation for a delta-distribution. It is used to test whether $R_{AA}$ measurements are sensitive to the shape of the jet-quenching or only the mean value for a given class of events.
@@ III $R_{\mathrm{AA}}$ for central Pb-Pb
As a further comparison, the distribution for the product of the $\alpha$ and $k$ parameters is shown in Figure 6, along with the correlation with $\beta$ for both the diagonal and covariant systematic errors. This product is proportional to the $p_{T}$-independent component of the mean $p_{T}$-loss, and the fact that it is well-constrained is consistent with the conclusion that only the mean of the $p_{T}$-dependent jet-quenching is required to adequately describe measurements of jet-$R_{AA}$.
@@ V Summary and Conclusions
We have constructed a simple, empirical model for jet-quenching in heavy-ion collisions. This model uses a 2-parameter product of a power-law and logarithm in the transverse momentum to describe the mean energy-loss. This model was extended by adding an additional parameter to account for a gamma-distribution $p_{T}$, but this extension performed no better than a delta-function in a Bayesian fit to the $\sqrt{s_{NN}}=5.02$ TeV 0-10% central Pb-Pb $R_{AA}$ measured by ATLAS. We conclude from this study that $R_{AA}$ measurements constrain only the mean value of $\Delta p_{T}$ and have little sensitivity to the shape of the $\Delta p_{T}$ distribution.

### v2

@@ I Introduction
In this work we focus primarily on the centrality and $p_{T}$ dependence of the jet nuclear-modification factor, $R_{AA}$, measured for Pb+Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV, as measured by the ATLAS Collaboration using the anti-$k_{T}$ algorithm with jet radius, $R=0.4$ [23]. These data were selected for the wide range in centrality and $p_{T}$ coverage and quantitative assessment of each of the systematic error components. Final fit results will also be compared to similar measurements by the CMS and ALICE collaborations [24, 25] collaborations. The simple model employed is an extension of the approach adopted in [20], in which the jet-energy loss distribution is described by a $p_{T}$ dependent function and applied to p-p jet cross-section to fit central $R_{AA}$ values using Bayesian methods. This work is also similar to [26], which also employs a Bayesian model to simultaneously fit both $R_{AA}$ and the di-jet assymetry, $x_{J}$, for central PbPb collisions to constrain jet-quenching for quark and gluon jets separately, but in this work we also extend the simple model to describe the centrality dependence of $R_{AA}$ by incorporating geometric information provided by the TRENTo  model [16]. The parameters are refit to the centrality-dependent $R_{AA}$ values under two different assumptions for the correlation of systematic errors. The jet-quenching distributions of the model are described in Sec. II, and Sec. III provides comparisons to central $R_{AA}$ measurements. Centrality dependent extensions and comparisons are provided in Sec. III, and the summary and conclusions are given in Sec. V. The treatment of systematic error correlations and impacts on fits are explored further in the Appendices. In Sec. A we compare different approaches for constructing the covariance error matrix, and the impacts of the different approaches are demonstrated in Sec. B.
@@ II Jet-quenching Distributions
The first step in developing a simple model for energy loss is to parameterize unquenched, inclusive jet cross-sections in $\sqrt{s_{NN}}=5.02$ TeV proton-proton collisions. For this we use the ATLAS inclusive jet cross-section measurements in the rapidity range $|y|<2.1$ [27], and fit to Eq. 1 for values of $800100<p_{T}<$ 800 GeV, sufficient to cover the ATLAS jet-$R_{AA}$ measurements spanning $631158<p_{T}<$ 631 GeV,
$\frac{d\sigma^{2}_{\rm jet}}{dp_{T}dy}=\frac{(c_{1}/p_{T})^{[c_{2}+c_{3}\log(p_{T})]}}{1+\exp[(p_{T}-c_{4})/c_{5}]}.$ (1)
@@
As noted in [26], the gamma-distribution is motivated by its equivalence to a convolution of $k$ exponentials, thereby providing a quasi-physical interpretation as the number of parton scatterings. There is no physical interpretation for a delta-distribution. It is used to test whether $R_{AA}$ measurements are sensitive to the shape of the jet-quenching or only the mean value for a given class of events.
@@ III $R_{\mathrm{AA}}$ for central Pb-Pb
As a further comparison, the distribution for the product of the $\alpha$ and $k$ parameters is shown in Figure 6, along with the correlation with $\beta$ for both the diagonal and covariant systematic errors. This product is proportional to the $p_{T}$-independent component of the mean $p_{T}$-loss, and the fact that it is well-constrained is consistent with the conclusion that only the mean of the $p_{T}$-dependent jet-quenching is required to adequately describe measurements of jet-$R_{AA}$. A gamma-distribution for jet-quenching can be properly constrained by also including the dijet asymmetry, $x_{J}$, as shown in  [26].
@@ V Summary and Conclusions
We have constructed a simple, empirical model for jet-quenching in heavy-ion collisions. This model uses a 2-parameter product of a power-law and logarithm in the transverse momentum to describe the mean energy-loss. The model was extended by adding an additional parameter to account for a gamma-distribution $p_{T}$, but this extension performed no better than a delta-function in a Bayesian fit to the $\sqrt{s_{NN}}=5.02$ TeV 0-10% central Pb-Pb $R_{AA}$ measured by ATLAS. We conclude from this study that $R_{AA}$ measurements constrain only the mean value of $\Delta p_{T}$ and have little sensitivity to the shape of the $\Delta p_{T}$ distribution. Extending the model to compare to the dijet asymmetry may be needed to properly constrain the gamma-distribution, as shown in  [26].
@@
The python code for this model is publicly available [31], and can be used to generate all figures in this work.